OpenStudy (anonymous):

I know how to find slope but idk how to figure this out. someone help! The equation of line AB is y = 5x + 1. Write an equation of a line parallel to line AB in slope-intercept form that contains point (4, 5). y = 5x − 15 y = 5x + 15 y = 1 over 5x + 21 over 5 y = 1 over 5x − 29 over 5 I know how to find slope but idk how to figure this out. someone help!

OpenStudy (texaschic101):

y = 5x + 1 The slope of this line is 5. A parallel line will have the same slope. Now use y = mx + b, and sub in our slope(m) = 5 and our points (4,5) into the formula and solve for b (the y intercept) y = mx + b 5 = 4(5) + b 5 = 20 + b 5 - 20 = b - 15 = b Therefore, our parallel line is : y = 5x - 15

OpenStudy (anonymous):

thank you so much !! it means a lot

OpenStudy (texaschic101):

do you understand everything or do you have any questions ?

OpenStudy (anonymous):

no thank you, i understand the lesson now (:

OpenStudy (texaschic101):

good to hear :) glad I could help